OpenStudy (anonymous):

slope of a line whose equation is 4x-6y=5

OpenStudy (anonymous):

use the slope formula

OpenStudy (texaschic101):

put your equation in y = mx + b form, and m is your slope

OpenStudy (texaschic101):

let me show you... 4x - 6y = 5 -- subtract 4x from both sides -6y = -4x + 5 -- divide both sides by -6 y = -4/-6x + 5/-6 y = 2/3x - 5/6 the number in the m position is 2/3. That is your slope

OpenStudy (anonymous):

Thank you! I am a student who has returned to school after 22 years....the math is...well thanks for the help!!